# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v2

@@ Abstract
Autonomous mobile robots must maintain safety, but should not sacrifice performance, leading to the classical reach-avoid problem. This paper seeks to compute trajectory plans for which a robot is guaranteed to reach a goal and to avoid obstacles in the specific near danger case that the obstacles and goal are near each other. The proposed method builds off of a common approach of using a simplified planning model to generate plans, which are then tracked using a high-fidelity tracking model and controller. Existing safe planning approaches use reachability analysis to overapproximate the error between these models, but this introduces additional numerical approximation error and thereby conservativeness that prevents goal reaching. The present work instead proposes a Piecewise Affine Reach-avoid Computation (PARC) method to tightly approximate the reachable set of the planning model. With PARC, the main source of conservativeness is the model mismatch, which can be mitigated by careful controller and planning model design. The utility of this method is demonstrated through extensive numerical experiments in which PARC outperforms state-of-the-art reach avoid methods in near-danger goal reaching. Furthermore, in a simulated demonstration, PARC enables generation of provably-safe extreme vehicle dynamics drift parking maneuvers.
@@ I Introduction
The increasing deployment of autonomous robots in daily tasks necessitates addressing complex operational challenges. Many of these tasks can be conceptualized as reach-avoid problems, wherein a robot must steer its states into a set of goal states without violating safety-critical constraints [1]. Achieving both liveness – the assurance of goal fulfillment [2] – and adherence to safety constraints is pivotal in these scenarios. In this paper, we are particularly interested in goal reaching near danger, meaning that an obstacle is near enough to a desired goal location to prevent typical robust planning or control approaches from reaching the goal. This often occurs when a goal and obstacles are closer to each other than one width of a robot’s body, such as for a quadrotor drone flying through a narrow gap or an autonomous car parallel parking.
@@
The first approach is control intervention, wherein a motion planner proposes actions, while a safety filter [3, 4] prevents the robot from executing actions for which its reachable set would intersect obstacles. This approach often requires compensating for the worst-case behavior of the motion planner [5, 6], which we find leads to conservativeness (i.e., sacrificing liveness) in our experiments (Section V). The second approach is a planner-tracker framework, wherein a motion planner proposes plans that are safe by leveraging the knowledge of a low-level controller’s tracking ability. This controller can be synthesized [7, 8] or hand-engineered [9, 10]; the key difference is that the planner need not be treated as adversarial, which reduces conservativeness. However, this approach typically limits the space of possible plans due to the need to compute the robot’s tracking ability. Furthermore, one must compensate for tracking error between the simplified planning model and the high-fidelity tracking model (i.e., the robot cannot perfectly track plans).
@@
Both control intervention and planner-tracker approaches suffer from challenges in numerical representation due to the need to underapproximate the set of states that reach the target set and overapproximate those that avoids obstacles when guaranteeing safety and liveness in implementation. Implementations such as sums-of-squares polynomials [7, 9, 11], grid-based PDE solutions [5], or polytope methods [10, 12] all introduce additional conservativeness purely due to compensating for numerical representation error. This challenge is exacerbated for systems with high-dimensional, nonlinear dynamics or observations; machine learning can address some of these concerns [13, 14, 15, 16, 17], at the expense of losing guarantees.
@@
In this work, we leverage two key insights to perform the goal-reaching near danger. First, as mentioned before, planner-tracker methods can reduce conservativeness by treating the planner and tracking controller as cooperative, as opposed to adversarial. Second, by representing our motion planner as a piecewise affine (PWA) system [18], we can tightly approximate its reachable set, with conservativeness mainly due to tracking controller and planning model design. We build off of a long history of discrete-time PWA system analysis in optimal control (e.g., [19, 20, 21]), which we extend to the continuous-time reach-avoid problem for robot motion planning. We demonstrate that this unlocks a faster and less conservative approach to a variety of reach-avoid problems from the robotics literature, and provides the solution to a reachable set computation for drift parking vehicle dynamics. An overview and example of our approach is shown in Fig. 1.
Contributions:
@@
    We propose a scheme to incorporate tracking error into PARC to underapproximate the reach set and overapproximate the avoid set for a trajectory planning model. This allows us to represent the BRAS that compactly represents all trajectory plans that reach the goal and avoid obstacles when tracked by the robot, represented by a high-fidelity dynamical model.
@@
    We stress-test PARC on examples drawn from the safe control and planning literature: a 4-D TurtleBot, a 10-D near-hover quadrotor, and a 13-D quadrotor. We show that PARC outperforms state-of-the-art methods such as FaSTrack [5], Neural CLBF [15], and RTD [9, 10]. We also demonstrate PARC on a simulated 6-D vehicle model to plan safe, extreme drift parallel parking maneuvers.
Paper Organization: Next, we introduce mathematical objects used throughout our formulation (Section II). We then present our problem formulation (Section III) and proposed approach (Section IV). We assess PARC with numerical experiments (Section V), then demonstrate a drift parking extreme vehicle dynamics example in simulation, extending beyond the state of the art in reachability-based safe motion-planning and control methods (Section VI). We discuss takeaways and limitations in Section VII. Note, we have extensive appendices, and call out key results in the text.
@@ II-B H-Polytopes
In this work, we represent all sets as H-polytopes or unions of H-polytopes, which have extensive algorithm and toolbox support [22, 23]. An ${n}$-dimensional H-polytope $\mathcal{P}(\mathbf{A},\mathbf{b})\subset\mathbb{R}^{{n}}$ is a convex set parameterized by ${n}_{\mathrm{\textnormal{h}}}$ linear constraints $\mathbf{A}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}}$ and $\mathbf{b}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}}$:
@@
Consider a pair of H-polytopes $P_{1}=\mathcal{P}(\mathbf{A}_{1},\mathbf{b}_{1})$, $P_{2}=\mathcal{P}(\mathbf{A}_{2},\mathbf{b}_{2})$. We make use of intersections $P_{1}\cap P_{2}$, Minkowski sums ($P_{1}\oplus P_{2}=\left\{\mathbf{x}+\mathbf{y}\ |\ \mathbf{x}\in P_{1},\mathbf{y}\in P_{2}\right\}$), Pontryagin differences ($P_{1}\ominus P_{2}=\left\{\mathbf{x}\in P_{1}\ |\ \mathbf{x}+\mathbf{y}\in P_{1}\ \forall\ \mathbf{y}\in P_{2}\right\}$), set differences $P_{1}\setminus P_{2}$, Cartesian products $P_{1}\times P_{2}$, and convex hulls $\mathrm{\textnormal{conv}}\!\left(P_{1},P_{2}\right)$. We also project an $n$-dimensional polytope $P$ into dimensions $p:q$ as $\mathrm{\textnormal{proj}}_{{p}:{q}}\!\left(P\right)$. These operations are detailed in Appendix -B.
@@ II-C PWA Systems
For all $\mathbf{x}(0)$ that share the same mode sequence, the dynamics reduce from time-variant PWA to time-variant affine up to time $t_{\mathrm{\textnormal{f}}}$. This is the core idea that enables the efficient underapproximating BRS computation of PARC. Note, taking advantage of fixed mode sequences is well-established in robotics and controls (e.g., [19, 24, 20, 25]), but we believe we are the first to show its utility in safe reach-avoid computation of robot planning models.
@@
When ${n}_{\mathrm{\textnormal{PWA}},t}=1$ for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, one obtains a discrete-time, time-variant affine system as a particular case of this problem formulation.
@@ II-D One Step BRS via Inverse Affine Map
Finally, we take advantage of the H-polytope analytic solution for inverse affine maps to enable BRS computation using the following well-known lemma (c.f., [19, 26, 22]).
@@
We take advantage of the fact that repeated applications of Lemma 1 to an H-polytope still produce an H-polytope. Notice that (8) does not require invertibility of $\mathbf{C}$, conferring an advantage over other set polytope representations such as constrained zonotopes [27]. This makes H-polytopes an excellent choice for representing the BRS.
@@ III-A Planning Model
A planning model provides a low-fidelity but fast-to-compute model of the robot’s dynamics. In this work, we use parameterized reference trajectories, which can allow quick computation and strong safety guarantees [9, 10, 28].
@@ III-B Tracking Model
where $\mathbf{z}_{0}\in Z\subset\mathbb{R}^{{n}_{z}}$ is the initial state. The tracking state $\mathbf{z}$ includes the planning state $\mathbf{p}$, i.e., $\mathbf{z}=[\mathbf{p}^{\intercal},\mathbf{z}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$. Given a plan $\mathbf{p}$ parameterized by $\mathbf{k}$, we use a feedback tracking controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}(t,\mathbf{z}(t);\mathbf{k})\in U$ (e.g., PID or MPC) to control the tracking model. We assume $Z$ and $U$ are compact and the dynamics are Lipschitz such that trajectories exist [29]. For a particular $\mathbf{z}_{0}$, we denote the resulting, unique realized trajectory (i.e., solution to (13)) as $\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})$ at time $t$.
@@
Designing the feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ is not the focus of this work. Conveniently, specifying a parameterized (i.e., restricted) space of plans $\mathbf{p}$ simplifies controller design. Our overall goal is to compensate for the fact that the controller is imperfect.
@@ III-C Reach-Avoid Problem
Note, we only consider static obstacles, because our goal is to plan a robot’s motion near danger. The danger resulting from moving or adversarial obstacles depends upon accurately predicting obstacle motion, which introduces an independent variable unrelated to our purpose. We leave extension to dynamic scenes for future work; we anticipate we can adapt reachability-based methods such as [30].
@@
Remark 5.
A fixed time horizon is reasonable because we seek to compute a trajectory near danger; the robot is already near its goal, but reaching the goal is difficult. The BRAS is a large set far from obstacles that can be used as a target set for other methods to reach safely.
Remark 6.

### v3

@@ Abstract
Autonomous mobile robots must maintain safety, but should not sacrifice performance, leading to the classical reach-avoid problem. This paper seeks to compute trajectory plans for which a robot is guaranteed to reach a goal and to avoid obstacles in the specific near danger case, also known as a narrow gap, where the agent starts near the goal, but must navigate through tight obstacles that block its path. The proposed method builds off of a common approach of using a simplified planning model to generate plans, which are then tracked using a high-fidelity tracking model and controller. Existing safe planning approaches use reachability analysis to overapproximate the error between these models, but this introduces additional numerical approximation error and thereby conservativeness that prevents goal reaching. The present work instead proposes a Piecewise Affine Reach-avoid Computation (PARC) method to tightly approximate the reachable set of the planning model. PARC significantly reduces conservativeness through a careful choice of the planning model and set representation, along with an effective approach to handling time-varying tracking errors. The utility of this method is demonstrated through extensive numerical experiments in which PARC outperforms state-of-the-art reach avoid methods in near-danger goal reaching. Furthermore, in a simulated demonstration, PARC enables generation of provably-safe extreme vehicle dynamics drift parking maneuvers. A preliminary hardware demo on a TurtleBot3 also validates the method.
@@ I Introduction
The increasing deployment of autonomous robots in daily tasks necessitates addressing complex operational challenges. Many of these tasks can be conceptualized as reach-avoid problems, wherein a robot must steer its states into a set of goal states without violating safety-critical constraints [fisac2015reach]. Achieving both liveness – the assurance of goal fulfillment [alpern1985defining] – and adherence to safety constraints is pivotal in these scenarios. In this paper, we are particularly interested in collision-free goal reaching near danger, meaning that the agent starts near the goal, but must navigate through tight obstacles—typically within the width of a robot’s body—that block its path. As we will show, this problem setup is very difficult for existing robust planning and control approaches to provide reach-avoid guarantees on general dynamical systems due to over-conservativeness. We note that the specific “near danger” condition where these planners and controllers fail is task-dependent. Some examples include flying a quadrotor drone through a narrow gap, or parallel parking an autonomous car.
I-A Reach-Avoid Approaches and Challenges
@@
The first approach is control intervention, wherein a motion planner proposes actions, while a safety filter [wabersich2023data, hsu2023safety] prevents the robot from executing actions for which its reachable set would intersect obstacles. This approach often requires compensating for the worst-case behavior of the motion planner [chen2021fastrack, garg2021multi], which we find leads to conservativeness (i.e., sacrificing liveness) in our experiments (Section V). The second approach is a planner-tracker framework, wherein a motion planner proposes plans that are safe by leveraging the knowledge of a low-level controller’s tracking ability. This controller can be synthesized [majumdar2017funnel, tedrake2010lqr] or hand-engineered [kousik2020bridging, kousik2019safe]; the key difference is that the planner need not be treated as adversarial, which reduces conservativeness. However, this approach typically limits the space of possible plans due to the need to compute the robot’s tracking ability. Furthermore, one must compensate for tracking error between the simplified planning model and the high-fidelity tracking model (i.e., the robot cannot perfectly track plans).
@@
Both control intervention and planner-tracker approaches suffer from challenges in numerical representation due to the need to underapproximate the set of states that reach the target set and overapproximate those that avoids obstacles when guaranteeing safety and liveness in implementation. Implementations such as sums-of-squares polynomials [majumdar2017funnel, kousik2020bridging, landry2018reach], grid-based PDE solutions [chen2021fastrack], or polytope methods [kousik2019safe, althoff2021set] all introduce additional conservativeness purely due to compensating for numerical representation error. This challenge is exacerbated for systems with high-dimensional, nonlinear dynamics or observations; machine learning can address some of these concerns [bansal2021deepreach, michaux2023reachability, dawson2022safe, xiao2023barriernet, selim2022safe], at the expense of losing guarantees.
I-B Proposed Method Summary and Context
@@
In this work, we leverage two key insights to perform goal-reaching near danger. First, we follow the literature [kousik2020bridging, kousik2019safe] in using a planner-tracker framework to reduce conservativeness by treating the planner and tracking controller as cooperative, as opposed to adversarial. Second, by representing our motion planner as a piecewise affine (PWA) system [christophersen2007optimal], we tightly approximate its reachable set, with minimal conservativeness from numerical approximation error. We demonstrate how this unlocks an efficient and less conservative approach to a variety of reach-avoid problems from the robotics literature, and provides the solution to a reachable set computation for drift parking vehicle dynamics. An overview and example of our approach is shown in Fig. 1.
Our method fits in the context of both control and planning. We build off of a long history of discrete-time PWA system analysis in optimal control (e.g., [thomas2006robust, kerrigan2002optimal, rakovic2005robust]), which we extend to the continuous-time reach-avoid problem for robot motion planning. Furthermore, our approach is complementary to sampling-based motion planners such as RRT [lavalle1998rapidly]; our approach could be applied in the future to compute a safety funnel around a sampled motion plan, or we could leverage sampling to seed our reachability analysis.
I-C Contributions
@@
    We propose a scheme to incorporate tracking error into PARC to compute a subset of the reach set and overapproximate the avoid set for a trajectory planning model. This allows us to represent the BRAS that compactly represents all trajectory plans that reach the goal and avoid obstacles when tracked by the robot, represented by a high-fidelity dynamical model.
@@
    We stress-test PARC on examples drawn from the safe control and planning literature: a 4-D TurtleBot, a 6-D planar quadrotor, a 10-D near-hover quadrotor, and a 13-D general quadrotor. We show that PARC outperforms state-of-the-art methods such as FaSTrack [chen2021fastrack], Neural CLBF [dawson2022safe], RTD [kousik2020bridging, kousik2019safe], and KDF [verginis2022kdf] in near danger reach-avoid tasks due to reduced conservativeness from the combination of planning model and set representation choice and better implementation of tracking error. We also demonstrate PARC on a simulated 6-D vehicle model to plan safe, extreme drift parallel parking maneuvers, and on a 4-D physical TurtleBot as a preliminary hardware experiment.
Paper Organization: Next, we introduce mathematical objects used throughout our formulation (Section II). We then present our problem formulation (Section III) and proposed approach (Section IV). We assess PARC with numerical experiments (Section V), then demonstrate a drift parking extreme vehicle dynamics example in simulation, extending beyond the state of the art in reachability-based safe motion-planning and control methods (Section VI). We also showcase PARC’s ability to extend to real robots with a preliminary hardware experiment (Section VII). Finally, we discuss takeaways and limitations in Section VIII. Note, we have extensive appendices, and call out key results in the text.
@@ II-B H-Polytopes
In this work, we represent all sets as H-polytopes or unions of H-polytopes, which have extensive algorithm and toolbox support [herceg2013multi, althoff2015introduction]. An ${n}$-dimensional H-polytope $\mathcal{P}(\mathbf{A},\mathbf{b})\subset\mathbb{R}^{{n}}$ is a convex set parameterized by ${n}_{\mathrm{\textnormal{h}}}$ linear constraints $\mathbf{A}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}}$ and $\mathbf{b}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}}$:
@@
Consider a pair of H-polytopes $P_{1}=\mathcal{P}(\mathbf{A}_{1},\mathbf{b}_{1})$, $P_{2}=\mathcal{P}(\mathbf{A}_{2},\mathbf{b}_{2})$. We make use of intersections $P_{1}\cap P_{2}$, Minkowski sums ($P_{1}\oplus P_{2}=\left\{\mathbf{x}+\mathbf{y}\ |\ \mathbf{x}\in P_{1},\mathbf{y}\in P_{2}\right\}$), Pontryagin differences ($P_{1}\ominus P_{2}=\left\{\mathbf{x}\in P_{1}\ |\ \mathbf{x}+\mathbf{y}\in P_{1}\ \forall\ \mathbf{y}\in P_{2}\right\}$), set differences $P_{1}\setminus P_{2}$, Cartesian products $P_{1}\times P_{2}$, and convex hulls $\mathrm{\textnormal{conv}}\!\left(P_{1},P_{2}\right)$. We also project an $n$-dimensional polytope $P$ into dimensions $p:q$ as $\mathrm{\textnormal{proj}}_{{p}:{q}}\!\left(P\right)$, and slice it in the dimensions $p:q$ with respect to some constant $\mathbf{x}_{0}\in\mathbb{R}^{q-p+1}$ as $\mathrm{\textnormal{slice}}_{{p}:{q}}\!\left(P,\mathbf{x}_{0}\right)$. These operations are detailed in Appendix -B.
@@ II-C PWA Systems
For all $\mathbf{x}(0)$ that share the same mode sequence, the dynamics reduce from time-variant PWA to time-variant affine up to time $t_{\mathrm{\textnormal{f}}}$. Mathematically, the discrete-time, time-variant affine dynamics with respect to the mode sequence (4) is:
$\begin{split}&\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\\ &\forall\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{s_{t},t},\mathbf{b}_{s_{t},t}),\end{split}$ (5)
where $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
The reduction of a PWA system to a specific affine dynamics is the core idea that enables the efficient underapproximating BRS computation of PARC. Note, taking advantage of fixed mode sequences is well-established in robotics and controls (e.g., [thomas2006robust, desimini2020robust, kerrigan2002optimal, wensing2023optimization]), but we believe we are the first to show its utility in safe reach-avoid computation of robot planning models.
@@
As a particular case of this problem formulation, when ${n}_{\mathrm{\textnormal{PWA}},t}=1$ for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, one obtains the discrete-time, time-variant affine system (LABEL:eq:time_variant_affine) from the PWA system (2) and (3) automatically, as $S=(1,\cdots,1)$ always.
@@
As will be shown in Section IV-C, the use of linear interpolation to bridge between discrete and continuous time is necessary to enable avoid set computation with H-polytopes. Similar reachability techniques to ours have been shown to work with Bernstein polynomials [michaux2023can], but only for forward reachability, to the best of our knowledge. Nonetheless, it may still be possible for PARC to handle alternative interpolation methods such as cubic splines or Bézier curves by using other set representations such as polynomial zonotopes [kochdumper2020sparse]. In this paper, we opted for the most straightforward approach of linear interpolation and leave the investigation with other interpolation methods as future work.
@@ II-D One Step BRS via Inverse Affine Map
Finally, we take advantage of the H-polytope analytic solution for inverse affine maps to enable BRS computation using the following well-known lemma (c.f., [thomas2006robust, vincent2021reachable, herceg2013multi]).
@@
We take advantage of the fact that repeated applications of Lemma 1 to an H-polytope still produce an H-polytope. Notice that (9) does not require invertibility of $\mathbf{C}$, conferring an advantage over other set polytope representations such as constrained zonotopes [scott2016constrained]. This makes H-polytopes an excellent choice for representing the BRS.
@@ III-A Planning Model
A planning model provides a low-fidelity but easy-to-compute model of the robot’s dynamics. In this work, we use parameterized reference trajectories, which can allow quick computation and strong safety guarantees [kousik2020bridging, kousik2019safe, hess2014formal].
@@ III-B Tracking Model
where $\mathbf{z}_{0}\in Z\subset\mathbb{R}^{{n}_{z}}$ is the initial state. The tracking state $\mathbf{z}$ includes the planning state $\mathbf{p}$, i.e., $\mathbf{z}=[\mathbf{p}^{\intercal},\mathbf{z}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$. Given a plan $\mathbf{p}$ parameterized by $\mathbf{k}$, we use a feedback tracking controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}(t,\mathbf{z}(t);\mathbf{k})\in U$ (e.g., PID or MPC) to control the tracking model. We assume $Z$ and $U$ are compact and the dynamics are Lipschitz such that trajectories exist [grant2014theory]. For a particular $\mathbf{z}_{0}$, we denote the resulting, unique realized trajectory (i.e., solution to (14)) as $\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})$ at time $t$.
@@
Designing the feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ is not the focus of this work. Conveniently, specifying a parameterized (i.e., restricted) space of plans $\mathbf{p}$ simplifies controller design, as we find experimentally in Section V-D. Our overall goal is to compensate for the fact that the controller is imperfect.
@@ III-C Reach-Avoid Problem
Note, we only consider static obstacles, because our goal is to plan a robot’s motion near danger. The danger resulting from moving or adversarial obstacles depends upon accurately predicting obstacle motion, which introduces an independent variable unrelated to our purpose. We leave extension to dynamic scenes for future work, but we anticipate we can adapt reachability-based methods such as [vaskov2019towards, vaskov2019not].
@@
Several remarks are in order to discuss our fixed time horizon, choice of goals and obstacles, and choice of tracking error model.
Remark 5 (Fixed Final Time).
A fixed time horizon is reasonable because we seek to compute a trajectory near danger; the robot is already near its goal, but reaching the goal is difficult. The BRAS is ideally a large set of initial conditions far from obstacles that can be used as a target set for other methods to reach safely.
Remark 6 (Goal and Obstacles).
@@
Remark 7 (Disturbances to the System).
To simplify exposition, we treat tracking error as a lumped uncertainty representation between planning and control, encompassing multiple error sources – disturbances, state estimation, object detection, and trajectory prediction. Separately analyzing these sources would introduce confounding variables, complicating the interpretation of our results. Instead, we consolidate them into a single metric of tracking error. For example, one can design a robust tracking controller that yields a tracking error bound, robust to external disturbances and model uncertainties [chen2021fastrack, bechlioulis2014low]. Our planning level only considers the result of those errors as a tracking error bound without separately accounting for each. Alternatively, if the bounds of inertial uncertainties [michaux2023can], dynamic disturbance [selim2022safe, alanwar2023data], and perception uncertainty [kousik2020bridging] are known (which are strong assumptions), then planner-tracker frameworks similar to ours have been shown to maintain reach-avoid guarantees in the past by “buffering” the bounds into the tracking error [kousik2020bridging, vaskov2019towards, selim2022safe, chen2021fastrack]. In Section VII, we show the preliminary results of a hardware experiment that successfully includes localization drift from the inertial measurement unit (IMU) and the encoders.